OpenStudy (anonymous):

Consider the following functions: f(x) = 1 + 1 -- x g(x) = 1 -- x Find the following and give the domain of each: f + g f - g f x g f / g Domain = ALL REAL NUMBERS - { ? } ( for all )

OpenStudy (anonymous):

0

OpenStudy (blacksteel):

I'm assuming your functions are \[f = 1 + 1/x, g = 1/x\] Then f + g = 1 + 2/x. Since x is in the denominator, the domain must exclude 0 (since division by 0 is undefined). Then f+g has domain R - {0}. f-g = 1, so the domain is just R. \[f*g = 1/x + 1/x^2\] so again we have to exclude any x values that make the denominator 0, in this case x = 0, so the domain is R - {0}. \[f/g = (1 + 1/x)/(1/x) = (1+1/x)*x = x + 1\] This again has domain R.

OpenStudy (anonymous):

thank you, and thanks for breaking it down too

OpenStudy (blacksteel):

Any time!